OpenStudy (anonymous):

If the radius and height of a cylinder are multiplied by 3, the lateral area of the cylinder is multiplied by

OpenStudy (zepp):

If the radius and the height are multiplied by 3, that means that k = 3 We are looking for the area, which is k^2 k = 3 k^2 = 3^2 = 9